OpenStudy (anya_marie):

ok, so I'm doing my math homework and a random question about log comes up and do not remember anything about that... it wants me to evaluate log base 2 1/32... please help

OpenStudy (mathmale):

\[\log_{2} \frac{ 1 }{ 32 }=?\] Please express 32 as a power of 2, OK?

OpenStudy (mathmale):

2^2=4 2^3=8 and so on

OpenStudy (anya_marie):

okay... what do i do now?

OpenStudy (mathmale):

express 32 as a power of 2.

OpenStudy (anya_marie):

\[\log_{2} 1/32\] like this?

OpenStudy (kayders1997):

Now that's what they gave you alrady

OpenStudy (kayders1997):

Already

OpenStudy (anya_marie):

oh nevermind, okay i get it now... thank you!

OpenStudy (kayders1997):

No (typing errors)

OpenStudy (mathmale):

No. 2^1=2 2^2=4 2^3=8 Please express 32 as a power of 2 following these examples.

OpenStudy (kayders1997):

What is the answer?

OpenStudy (kayders1997):

You got

OpenStudy (mathmale):

2 to what power produces 32... ?

OpenStudy (kayders1997):

Or 2^x=32 what is x?

OpenStudy (anya_marie):

it doesn't come out even... i got \[4\sqrt{2}\]

OpenStudy (kayders1997):

Hmmmm how did you get that?

OpenStudy (anya_marie):

square root 32, do a tree because the calculator gave me a decimal, and simplify

OpenStudy (kayders1997):

Your nothing taking square root of 32

OpenStudy (anya_marie):

then how would you get x?

OpenStudy (kayders1997):

when you do this'll you have the base of 2 so that's going to be multiplied some amount of times to get 1/32

OpenStudy (mathmale):

Please follow the pattern I've typed out for you. 2^1 is "2 to the first power;" the answer is 2. 2^2 is "2 to the 2nd power;" the answer is 4. Please continue this patter until you discover which power of 2 gives you 32.

OpenStudy (anya_marie):

so it's \[1/32\div2?\]

OpenStudy (kayders1997):

No

OpenStudy (anya_marie):

its 5!

OpenStudy (mathmale):

Do NOT use a calculator for this. You need to be able to do these particular calculations in your head. What does 2^3 mean: Explain it in words, please. Can you explain why you're not comfortable doing this?

OpenStudy (kayders1997):

Because we want 2 to be multiplied so many times so 2 times 2 times 2 times 2 until we get to 32 just think of it that way for now

OpenStudy (mathmale):

Example: "The fourth power of 2 is 16." Your turn: "The fifth power of 2 is ...... ?

OpenStudy (anya_marie):

32

OpenStudy (mathmale):

"The 6th power of 2 is 64."

OpenStudy (kayders1997):

Yes so 2^5 power is 32 now how do you think you would get 1/32?

OpenStudy (mathmale):

Yes. 32=2^5. Now, rewrite 1/32 with a power of 2 in the denominator.

OpenStudy (anya_marie):

-5?

OpenStudy (kayders1997):

Yes

OpenStudy (mathmale):

Kayders asked the right question and y ou have given the right answer. Can you now finish the problem?

OpenStudy (anya_marie):

yeah... makes more sense. thank you both!

OpenStudy (mathmale):

What is the log to the base 2 of 1/32? What is the log to the base 2 of \[\frac{ 1 }{ 32 }?\]

OpenStudy (kayders1997):

:) glad to help

OpenStudy (anya_marie):

-5

OpenStudy (mathmale):

What is the log to the base 2 of \[\frac{ 1 }{ 2^5 }?\]

OpenStudy (mathmale):

Kayders?

OpenStudy (kayders1997):

What?

OpenStudy (mathmale):

Please give Anya_Marie feedback on her lastest answer.

OpenStudy (kayders1997):

Well it's right, I don't know what else to say 2^-5=1/32 remember if you have a fraction you will always have a negative power

OpenStudy (mathmale):

A_M: Thank you for sticking with us. sorry it took a while....but it looks as tho' you are understanding much better what to do to "solve" this problem. Thank you, kayders, for your input.

OpenStudy (mathmale):

yes, the correct answer is -5. That's the log to the base 2 of 1/32. Good.

OpenStudy (anya_marie):

it's okay. and thanks for helping... math really isn't my strongest point... and this was the last question on the homework and we haven't studied log in like 3 sections. kinda caught me off guard